A MOTHER IS 21 YEARS OLDER THAN HER CHILD. IN EXACTLY 6 YEARS THE MOTHER WILL BE 5 TIMES AS OLD AS THE CILD. WHAT IS THE CHILD,S AGE?
Answer by nyc_function(2741)

child's age now = x
mother's age now = x + 21
mother's age in 6 years =- (x + 21) + 6
child's age in 6 years = x + 6
Here is your equation:
(x + 21) + 6 = 5(x + 6)
Can you solve for x now?